OpenStudy (anonymous):

The number of hours, H, needed to manufacture X computer monitors is given by the linear function defined by H=kX+q, where k and q are constants. If it takes 270 hours to manufacture 100 computer monitors and 410 hours to manufacture 160 computer monitors, how many minutes are required to manufacture one computer monitor?

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

@zzr0ck3r

OpenStudy (zzr0ck3r):

ok so you have a linear equation y=mx+b where y = hours and x = computers, with two points (100,270), and (160,410) so the slope is (410-270)/160-100) = 140/60 = 7/3 so y=(7/3)x+b and we pick a point to find the y intercept, ill use (100,270) 270 = (7/3)(100)+b b = 270-(7/3)(100) = 110/3 so your equation has the form y = (7x+110)/3 we need to find the hours (y) to manufacture 1 computer (x=1) y = (7+110)/3 = 39 hours

Directrix (directrix):

I wish @sakigirl would start us off with a simpler question and then move up to the harder ones. We should rotate the responsibility for questions among ourselves. What do y'all think about that? @zzr0ck3r

OpenStudy (zzr0ck3r):

im fine with that, Its my anniversary and my wife is making dinner. But I will be here later for a while

Directrix (directrix):

What I mean is the responsibility for selecting and posting the questions, not necessarily answering them. Hey, the only thing more important than math is love so please celebrate your anniversary and don't worry about hurrying back here.

OpenStudy (anonymous):

@Directrix Well, these are just my homework questions

OpenStudy (anonymous):

I'm posting up the ones that I'm stuck on.

Directrix (directrix):

I knew that was something odd about them. They did not have the appearance of SAT questions.

Directrix (directrix):

@sakigirl Why are you doing homework during the summer, if it is summer where you are?

OpenStudy (anonymous):

@Directrix I attend SAT classes during the summer, sadly